# Redirection for Erasing Memory (REM): Towards a universal unlearning method for corrupted data

### v1

@@ Abstract
Machine unlearning is studied for a multitude of tasks, but specialization of unlearning methods to particular tasks has made their systematic comparison challenging. To address this issue, we propose a conceptual space to characterize diverse corrupted data unlearning tasks in vision classifiers. This space is described by two dimensions, the discovery rate (the fraction of the corrupted data that are known at unlearning time) and the statistical regularity of the corrupted data (from random exemplars to shared concepts). Methods proposed previously have been targeted at portions of this space and—we show—fail predictably outside these regions. We propose a novel method, Redirection for Erasing Memory (REM), whose key feature is that corrupted data are redirected to dedicated neurons introduced at unlearning time and then discarded or deactivated to suppress the influence of corrupted data. REM performs strongly across the space of tasks, in contrast to prior SOTA methods that fail outside the regions for which they were designed.
@@ 1 Introduction
Unlearning is the problem of removing the effect of a subset of training data from a trained model [21]. In this work, we consider a scenario where after having already trained a model on a dataset, we discover that a subset of the training data was accidentally mislabelled, low quality, or manipulated by an attacker, causing the model to make mistakes or exhibit unwanted behaviours. The goal of unlearning (UL) is to post-process the trained model to efficiently remove (the effect of) that corrupted data in order to restore the correct predictions and behaviours.
@@
While the problem of unlearning has attracted significant attention [29, 12, 9], and a plethora of methods have been proposed, there is still a lack of scientific understanding of the behaviours of these methods on different types of UL tasks, with only early work in this direction [32, 9]. This lack of a systematic understanding of when different methods fail or succeed is a fundamental blind spot as it hinders research progress and can lead to unpredictable failure in practice as shown in this work.
@@
To address this, our first contribution is a taxonomy of tasks for unlearning corrupted data shown in Fig. 1. Our taxonomy is based on the identification of two dimensions along which the behaviours of unlearning algorithms differ substantially. The first is the discovery rate, the fraction of corrupted data that have been identified and can be utilized by the unlearning algorithm to remove the effect of corruptions. Goel et. al [9] previously studied the effect of discovery rates on unlearning performance, finding that performance of algorithms developed for the full discovery case drops off suddenly as the discovery rate is lowered continuously. We take this study a step further by identifying a second dimension that affects performance significantly, both on its own and through interaction with the discovery rate. The second dimension is the statistical regularity of the corrupted data, tracing a spectrum from “spurious” corruptions (such as random mislabeling), towards structured corruptions that systematically affect related data points (such as a poison trigger that redirects all images on which it appears to a pre-specified class label). Our key finding is that, in our 2D space of tasks, different state-of-the-art (SOTA) methods succeed in different slices, but each fails predictably and catastrophically everywhere else. These methods are therefore risky to use in practice when the task specification is not known.
@@
Figure 1: We present a new taxonomy of unlearning tasks in terms of two dimensions: the regularity and the discovery rate of the corrupted data we wish to unlearn. The highlighted areas in (a) show tasks studied in prior work. Subplots (b-e) show an aggregate metric of unlearning performance (see Section 2) of different methods for different discovery rates >0% (x-axis) and regularities (y-axis), instantiated via the benchmarks of [9]. Better performance is shown in darker color. Prior methods succeed in only slices of this 2D space, mainly failing along the regularity axis.
@@
To address the need for a universal method for unlearning corrupted data that covers the 2D space, we present our second contribution: Redirection for Erasing Memory (REM), a new unlearning algorithm that is the first to perform strongly across our task space. REM employs a novel mechanism, inspired by the memorization mitigation method Example-Tied Dropout (ETD) [18], to redirect the corruption to a dedicated part of the network (newly initialized capacity added to the model by REM) that can then be conveniently dropped off or deactivated to unlearn even undiscovered corrupted data. We benchmark REM against SOTA unlearning methods as well as ETD on low, medium, and high regularity unlearning tasks for discovery rates ranging from 10% to 100%. Experimental results on CIFAR10 [16] and Street View House Numbers (SVHN) [20] with ResNet-9 [13] and Vision Transformer (ViT) [3] as well as different optimizers show that REM is the only method to perform well or equally well across the entire 2D space.
@@ 2 Problem Definition and Background
Machine unlearning applications range from privacy protection of individuals [19, 26, 12] to correcting errors in the model due to corrupted training data [8, 17, 9], each of which has different goals and evaluation metrics. In this work, we focus on unlearning corrupted data for classification tasks. Practical examples of corrupted data can range from simple data entry errors [27, 24] to large scale data poisoning attacks [1, 31].
@@
Problem definition. Let $\mathcal{D}_{\text{tr}}$ denote a training dataset, and $\mathcal{D}_{\text{test}}$ a held-out dataset of the same distribution. Further, let $\mathcal{A}$ denote a training algorithm, and $\theta=\mathcal{A}(\mathcal{D})$ the weights of a neural network that are learned by applying $\mathcal{A}$ on $\mathcal{D}$. We consider scenarios where some training data $\mathcal{D}_{\text{c}}\subset\mathcal{D}_{\text{tr}}$ is corrupted. Let $\mathcal{D}_{\text{f}}\subseteq\mathcal{D}_{\text{c}}$ denote the forget set comprised of corrupted data discovered after training. We refer to the rest of the training data, $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{r}}$ as the retain set. This formulation reduces to the most commonly-studied variation of unlearning when $\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{c}}$; i.e. the full set of data that we wish to unlearn is provided in the forget set. In that case, a straightforward but computationally expensive solution is to retrain from scratch to obtain a new model $\theta_{r}=\mathcal{A}(\mathcal{D}_{\text{r}})$ that is not affected by corrupted data. For “partial discovery”, defined as $\mathcal{D}_{\text{f}}\subset\mathcal{D}_{\text{c}}$, $\theta_{r}$ is not a solution since undiscovered corrupted data $\mathcal{D}_{\text{c}}\setminus\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{undiscovered}}$ is part of $\mathcal{D}_{\text{r}}$, so re-training on it does not eliminate the corruptions [9].
The problem we are interested in is to design a ($\mathcal{A}$, $\mathcal{U}$) pair of a learning algorithm $\mathcal{A}$ and post-processing (unlearning) algorithm $\mathcal{U}$ such that, when a forget set $\mathcal{D}_{\text{f}}$ is identified, $\mathcal{U}$ can efficiently post-process the trained model $\theta_{o}=\mathcal{A}(\mathcal{D}_{\text{tr}})$ to yield an unlearned model $\theta_{u}=\mathcal{U}(\theta_{o},\mathcal{D}_{\text{f}})$ that does not suffer from failure modes caused by $\mathcal{D}_{\text{c}}$, while having high utility. Note that some choices of $\mathcal{A}$ may facilitate the success of $\mathcal{U}$ to post-process for unlearning. We assume for simplicity that $\mathcal{U}$ can access $\mathcal{D}_{\text{tr}}$, as is common in the literature.
@@
Memorization and statistical regularities. Intuitively, [4] refers to an example $x$ as being “memorized” by a learning algorithm $\mathcal{A}$ in a training set $\mathcal{D}_{\text{tr}}$ if, in expectation, models obtained through recipe $\mathcal{A}(\mathcal{D}_{\text{tr}})$ are much more likely to make correct predictions on $x$ compared to models obtained through recipe $\mathcal{A}(\mathcal{D}_{\text{tr}}\setminus x)$. Intuitively, $x$ is memorized if it is required in $\mathcal{D}_{\text{tr}}$ to be correctly predicted. This will be the case for atypical / irregular examples [5], whereas typical / regular examples that are similar to many others in the dataset may be predicted correctly simply due to generalization (i.e. being learned through other samples), without having to be part of the training set. Similarly, [14] define a consistency score of an example $x$ to measure its statistical regularity via the expected accuracy on that example under models obtained by training on different subsets of $\mathcal{D}_{\text{tr}}$ that exclude $x$. Building upon this notion, we identify the statistical regularity of the corrupted data as a key factor influencing behaviours of unlearning algorithms.
@@
Example-Tied Dropout (ETD) is a training recipe proposed by [18] that separates the model parameters into those that intend to capture “generalizable information” that is shared between many data points, and those that intend to “memorize” information that is specific to one or a few data points. This separation is made within every layer of a neural network. During training, the computational paths of all data points pass through all of the generalization neurons (that is, the generalization neurons are never dropped out), but when it comes to the memorization partition, each example only passes through its dedicated path, based on an example-tied dropout mask that is randomly determined ahead of training and then fixed. This allows different pathways in the memorization partition to be “owned” by different examples, enabling those examples to encode their potential irregularities in their dedicated paths. On the other hand, example-specific irregularities won’t be as easily encoded in the generalization neurons, since those are shared by all examples. At inference time, the memorization neurons are dropped out, reducing or eliminating example-specific information. A similar approach has also been applied by [7] to prevent verbatim memorization in LLMs. We propose to leverage ETD as an UL algorithm, which can be seen as setting $\mathcal{A}$ to be ETD training followed by dropping out all memorization neurons post training, and setting $\mathcal{U}$ a no-op.
@@ 3 Related Work
Unlearning methods. We build upon the benchmarking selection of [9] to cover different state-of-the art approaches to unlearning and extend it with recent advances and overlooked baselines. [9] included Bad Teacher (BadT) [2], SCRUB [17], Selective Synaptic (SSD) [6], fine-tuning, and retraining from scratch. BadT randomizes labels on $\mathcal{D}_{\text{f}}$ by distilling from a randomly initialized network to induce forgetting and distilling the remainder of the samples $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}$. SCRUB alternates between a step of distillation away from the original model on $\mathcal{D}_{\text{f}}$ and a step towards the original model on $\mathcal{D}_{\text{r}}$ for model utility preservation on the remaining data. SSD determines disproportionately important parameters for $\mathcal{D}_{\text{f}}$ compared to $\mathcal{D}_{\text{tr}}$ in the model via the Fisher Information Matrix and dampens these parameters to induce unlearning. This leads to highly localized unlearning, only modifying $<2\%$ of model parameters. We replace SSD in our benchmarks with Potion [25] which builds upon SSD and is the SOTA in poison unlearning as defined by [9] (see Fig. 1). Potion iteratively increases the number of parameters in the model that it modifies analogously to SSD until unlearning of the poison trigger occurs. Potion stops unlearning as soon as it detects that the accuracy of the model on the forget set drops below a preset threshold, indicating that the poison has been unlearned. Fine-tuning continues to train the model on only $\mathcal{D}_{\text{r}}$, relying on catastrophic forgetting to remove $\mathcal{D}_{\text{f}}$. Retraining from scratch trains a new model using only $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}$. Ascent performs gradient ascent on $\mathcal{D}_{\text{f}}$.
@@
Investigation of unlearning limitations. [9] studies the effect of partial discovery on unlearning method effectiveness as shown in Fig. 1. For unlearning poisoned data, they demonstrate that all methods apart from [6] fail to remove the effect of the poison when only a subset of the poisoned examples is given (partial discovery). They also consider another setup, inter-class confusion, where samples from two classes are swapped with each other, where no method shows satisfactory performance in the partial discovery case. [32] identify interpretable characteristics of forget sets that substantially affect the difficulty of the problem, but they don’t consider the statistical regularity of the forget set in their investigation. [23] show that standard unlearning methods designed for erasing knowledge may fail at restoring knowledge present prior to corruption. This relates to our metrics of healing, where we are interested in redirection to the correct label, rather than just avoiding predicting the incorrect (e.g. poisoned) label. However, their investigation is in the context of LLMs and they don’t establish a distinction between tasks of different discovery rates or regularities.
4 Our Taxonomy of Unlearning Tasks and Limitations of Existing Methods
@@
Taxonomy. Our first contribution is to propose a new taxonomy of unlearning tasks according to two dimensions: the discovery rate and the statistical regularity of the corrupted data that we wish to unlearn. The discovery-rate dimension was previously discussed and empirically investigated in [9] but is insufficient to explain model failures, as methods such as Potion [25] excel at some low discovery tasks but then fail at other full-discovery tasks. We further identify regularity as a key dimension that affects the behaviour of unlearning algorithms both on its own and in conjunction with the discovery rate. Our key observation is that, as shown in Fig. 1, each prior method may excel on a slice of our 2D space of unlearning tasks but fails catastrophically everywhere else. We discuss the limitations of different algorithms below.
@@
Discovery rate. Unlearning methods that are designed for the traditional unlearning setting of full discovery fail catastrophically for lower discovery percentages as described by [9]. Retraining from scratch and unlearning methods that utilize the retain set, e.g. SCRUB, fail due to requiring a clean retain set for utility preservation, but in the partial discovery case, the retain set is contaminated and thus leads to reintroduction of corruptions (see Fig. 1 (b, e)). Methods such as Potion and SSD that only use $\mathcal{D}_{\text{f}}$ cause significant model utility damage due to the missing repair step on $\mathcal{D}_{\text{r}}$ [9, 6].
@@
Regularity refers to the self-similarity of the corrupted data $\mathcal{D}_{\text{c}}$. Informally, we say that a corruption has low regularity if the corrupted data does not share any structure or common patterns. For example, a random mislabelling of a randomly-chosen subset of examples is a low-regularity corruption because the self-similarity between samples that end up with the same label change is low. On the other hand, a poisoning attack that introduces a trigger to each image whose label it wants to redirect results in a high-regularity corruption since all corrupted images share the visual pattern of the poison trigger and are all labeled in the same way. In the context of the commonly studied problem of full discovery unlearning, class unlearning [10] and random subset unlearning [2] are high and low regularity tasks, respectively. Several measures could be used to formalize regularity. We advocate for the consistency score (C-score) of [14], that measures the expected accuracy of a model on held-out samples from the training set. Their precomputed CIFAR10 scores align with the regularity rankings of our tasks (see Fig. 13 in [14]). Adapted to $\mathcal{D}_{\text{c}}$, C-scores can be computed with $n=0,1,...|\mathcal{D}_{\text{c}}|$ for instances $x$ and labels $y$, where $\mathbb{E}^{r}$ is empirical averaging over $r$ samples [14]:
$\hat{C}_{\hat{\mathcal{D}_{\text{c}}}}(x,y)=\mathbb{E}_{n}[\hat{\mathbb{E}}^{r}_{D\stackrel{{\scriptstyle n}}{{\sim}}\mathcal{D}_{\text{c}}}\left[\mathbb{P}(f(x;D\backslash\{(x,y)\})=y)\right]].$ (1)
@@
Experimental setup. We evaluate tasks of low, medium, and high regularity with ten discovery rates for each task (10%-100%). Low regularity is represented by random label swaps of $n$ samples to an incorrect label [18]. In this setting, there is no regularity linking the samples or the new labels together. Medium regularity uses the inter-class confusion setup of [8] where $n$ samples of two classes are swapped with each other. The highest regularity task uses a poison trigger to redirect to class 0 as defined by [9]. We measure the accuracy of $\theta_{u}$ on two types of data: (i) corrupted data (using the clean labels to compute accuracy), and (ii) non-corrupted data. The former is used to measure “healing”, i.e. whether the label prediction has been “redirected” to the correct label, while the latter is used to measure “utility”, i.e. ability to predict correctly on non-corrupted data. We perform experiments on an A100 (40GB) GPU with ResNet-9 [13] and Vision Transformer (ViT) [3] models, stochastic gradient descent (SGD) and Adam [15] optimizers, and CIFAR10 [16] and Street View House Numbers (SVHN) [20] to show generalizability across models, optimizers and datasets.

### v2

@@ Abstract
Machine unlearning is studied for a multitude of tasks, but specialization of unlearning methods to particular tasks has made their systematic comparison challenging. To address this issue, we propose a conceptual space to characterize diverse corrupted data unlearning tasks in vision classifiers. This space is described by two dimensions, the discovery rate (the fraction of the corrupted data that are known at unlearning time) and the statistical regularity of the corrupted data (from random exemplars to shared concepts). Methods proposed previously have been targeted at portions of this space and, as we show, fail predictably outside these regions. We propose Redirection for Erasing Memory (REM), whose key feature is that corrupted data are redirected to dedicated neurons introduced at unlearning time and then discarded or deactivated to suppress the influence of corrupted data. REM performs strongly across the space of tasks, in contrast to prior SOTA methods that fail outside the regions for which they were designed.
@@ 1 Introduction
Unlearning is the problem of removing the effect of a subset of training data from a trained model (Nguyen et al., 2022). In this work, we consider a scenario where after having already trained a model on a dataset, we discover that a subset of the training data was accidentally mislabelled, low quality, or manipulated by an attacker, causing the model to make mistakes or exhibit unwanted behaviours. The goal of unlearning (UL) is to post-process the trained model to efficiently remove (the effect of) that corrupted data in order to restore the correct predictions and behaviours.
@@
While the problem of unlearning has attracted significant attention (Triantafillou et al., 2024; Hayes et al., 2024; Goel et al., 2024), and a plethora of methods have been proposed, there is still a lack of scientific understanding of the behaviours of these methods on different types of UL tasks, with only early work in this direction (Zhao et al., 2024; Goel et al., 2024). This lack of an understanding when established unlearning methods fail or succeed is a fundamental blind spot as it hinders research progress and can lead to unpredictable failure in practice, as shown in this work.
@@
To address this, our first contribution is a taxonomy of tasks for unlearning corrupted data shown in Fig. 1. Our taxonomy is based on the identification of two dimensions along which the behaviours of unlearning algorithms differ substantially. The first is the discovery rate, the fraction of corrupted data that have been identified and can be utilized by the unlearning algorithm to remove the effect of corruptions. Goel et al. (2024) previously studied the effect of discovery rates on unlearning performance, finding that performance of algorithms developed for the full discovery case drops off suddenly as the discovery rate is lowered continuously. We take this study a step further by identifying a second dimension that affects performance significantly, both on its own and through interaction with the discovery rate. The second dimension is the statistical regularity of the corrupted data, tracing a spectrum from “spurious” corruptions (such as random mislabeling), towards structured corruptions that systematically affect related data points (such as a poison trigger that redirects all images on which it appears to a pre-specified class label). Our key finding is that, in our 2D space of tasks, different state-of-the-art (SOTA) methods succeed in different slices, but each fails predictably and catastrophically everywhere else. These methods are therefore risky to use in practice when the task specification is not known.
@@
Figure 1: We present a new taxonomy of unlearning tasks in terms of two dimensions: the regularity and the discovery rate of the corrupted data we wish to unlearn. The highlighted areas in (a) show tasks studied in prior work. Subplots (b-e) show an aggregate metric of unlearning performance (see Section 2) of methods for different discovery rates >>0% (x-axis) and regularities (y-axis), instantiated via the benchmarks of (Goel et al., 2024). Darker color is better. Prior methods succeed only in slices of this 2D space, mainly failing along the regularity axis. Representative prior work: [A] Maini et al. (2023), [B] Schoepf et al. (2024b), [C] Chundawat et al. (2023); Kurmanji et al. (2023), [D] Goel et al. (2024), [E] Zhao et al. (2024); Foster et al. (2024); Chundawat et al. (2023); Kurmanji et al. (2023), [F] Chundawat et al. (2023); Kurmanji et al. (2023)
@@
To address the need for a universal method for unlearning corrupted data that covers the 2D space, we present our second contribution: Redirection for Erasing Memory (REM), a new unlearning algorithm that is the first to perform strongly across our task space. REM employs a novel mechanism, inspired by the memorization mitigation method Example-Tied Dropout (ETD) (Maini et al., 2023), to redirect the corruption to a dedicated part of the network (newly initialized capacity added to the model by REM) that can then be conveniently dropped off or deactivated to unlearn even undiscovered corrupted data of any regularity. We benchmark REM against SOTA unlearning methods as well as ETD on low, medium, and high regularity unlearning tasks for discovery rates ranging from 10% to 100%. Experimental results on CIFAR10 (Krizhevsky et al., 2009) and Street View House Numbers (SVHN) (Netzer et al., 2011) with ResNet-9 (He et al., 2016) and Vision Transformer (ViT) (Dosovitskiy et al., 2021) as well as different optimizers show that REM is the only method to perform well or equally well across the entire 2D space.
@@ 2 Problem Definition and Background
Machine unlearning applications range from privacy protection of individuals (Neel et al., 2021; Sekhari et al., 2021; Hayes et al., 2024) to correcting errors in the model due to corrupted training data (Goel et al., 2022; Kurmanji et al., 2023; Goel et al., 2024), each of which has different goals and evaluation metrics. In this work, we focus on unlearning corrupted data for classification tasks. Practical examples of corrupted data can range from simple data entry errors (Tanno et al., 2022; Schoepf et al., 2024a) to large scale data poisoning attacks (Carlini et al., 2024; Zhang et al., 2024b).
@@
Problem definition. Let $\mathcal{D}_{\text{tr}}$ denote a training dataset, and $\mathcal{D}_{\text{test}}$ a held-out dataset of the same distribution. Further, let $\mathcal{A}$ denote a training algorithm, and $\theta=\mathcal{A}(\mathcal{D})$ the weights of a neural network that are learned by applying $\mathcal{A}$ on $\mathcal{D}$. We consider scenarios where some training data $\mathcal{D}_{\text{c}}\subset\mathcal{D}_{\text{tr}}$ is corrupted. Let $\mathcal{D}_{\text{f}}\subseteq\mathcal{D}_{\text{c}}$ denote the forget set comprised of corrupted data discovered after training. We refer to the rest of the training data, $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{r}}$ as the retain set. This formulation reduces to the most commonly-studied variation of unlearning when $\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{c}}$; i.e. the full set of data that we wish to unlearn is provided in the forget set. In that case, a straightforward but computationally expensive solution is to retrain from scratch to obtain a new model $\theta_{r}=\mathcal{A}(\mathcal{D}_{\text{r}})$ that is not affected by corrupted data. For “partial discovery”, defined as $\mathcal{D}_{\text{f}}\subset\mathcal{D}_{\text{c}}$, $\theta_{r}$ is not a solution since undiscovered corrupted data $\mathcal{D}_{\text{c}}\setminus\mathcal{D}_{\text{f}}=\mathcal{D}_{\text{undiscovered}}$ is part of $\mathcal{D}_{\text{r}}$, so re-training on it does not eliminate the corruptions (Goel et al., 2024). The problem we are interested in is to design a ($\mathcal{A}$, $\mathcal{U}$) pair of a learning algorithm $\mathcal{A}$ and post-processing (unlearning) algorithm $\mathcal{U}$ such that, when a forget set $\mathcal{D}_{\text{f}}$ is identified, $\mathcal{U}$ can efficiently post-process the trained model $\theta_{o}=\mathcal{A}(\mathcal{D}_{\text{tr}})$ to yield an unlearned model $\theta_{u}=\mathcal{U}(\theta_{o},\mathcal{D}_{\text{f}})$ that does not suffer from failure modes caused by $\mathcal{D}_{\text{c}}$, while having high utility. Note that some choices of $\mathcal{A}$ may facilitate the success of $\mathcal{U}$ to post-process for unlearning. We assume for simplicity that $\mathcal{U}$ can access $\mathcal{D}_{\text{tr}}$, as is common in the literature.
@@
Memorization and statistical regularities. Feldman (2020) refers to an example $x$ as being “memorized” by a learning algorithm $\mathcal{A}$ in a training set $\mathcal{D}_{\text{tr}}$ if, in expectation, models obtained through recipe $\mathcal{A}(\mathcal{D}_{\text{tr}})$ are much more likely to make correct predictions on $x$ compared to models obtained through recipe $\mathcal{A}(\mathcal{D}_{\text{tr}}\setminus x)$. Intuitively, $x$ is memorized if it is required in $\mathcal{D}_{\text{tr}}$ to be correctly predicted. This will be the case for atypical / irregular examples (Feldman and Zhang, 2020), whereas typical / regular examples that are similar to many others in the dataset may be predicted correctly simply due to generalization (i.e. being learned through other samples), without having to be part of the training set. Similarly, Jiang et al. (2020) define a consistency score of an example $x$ to measure its statistical regularity via the expected accuracy on that example under models obtained by training on different subsets of $\mathcal{D}_{\text{tr}}$ that exclude $x$. Building upon this notion, we identify the statistical regularity of the corrupted data as a key factor influencing behaviours of unlearning algorithms.
@@
Example-Tied Dropout (ETD) is a training recipe proposed by Maini et al. (2023) that separates model parameters into those that intend to capture “generalizable information” that is shared between many data points, and those that intend to “memorize” information that is specific to one or a few data points. This separation is made within every layer of a neural network. During training, the computational paths of all data points pass through all of the generalization neurons (that is, the generalization neurons are never dropped out), but when it comes to the memorization partition, each example only passes through its dedicated path, based on an example-tied dropout mask that is randomly determined ahead of training and then fixed. This allows different pathways in the memorization partition to be “owned” by different examples, enabling those examples to encode their potential irregularities in their dedicated paths. On the other hand, example-specific irregularities won’t be as easily encoded in the generalization neurons, since those are shared by all examples. At inference time, the memorization neurons are dropped out, reducing or eliminating example-specific information. A similar approach has been applied by Ghosal et al. (2025) to prevent verbatim memorization in LLMs. We propose ETD as an UL algorithm, which can be seen as setting $\mathcal{A}$ as ETD training followed by dropping out all memorization neurons post training, and setting $\mathcal{U}$ a no-op.
@@ 3 Related Work
Unlearning methods. We build upon the benchmarking selection of Goel et al. (2024) to cover different state-of-the art approaches to unlearning and extend it with recent advances and overlooked baselines. Goel et al. (2024) included Bad Teacher (BadT) (Chundawat et al., 2023), SCRUB (Kurmanji et al., 2023), Selective Synaptic (SSD) (Foster et al., 2024), fine-tuning, and retraining from scratch. BadT randomizes labels on $\mathcal{D}_{\text{f}}$ by distilling from a randomly initialized network to induce forgetting and distilling the remainder of the samples $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}$. SCRUB alternates between a step of distillation away from the original model on $\mathcal{D}_{\text{f}}$ and a step towards the original model on $\mathcal{D}_{\text{r}}$ for model utility preservation on the remaining data. SSD determines disproportionately important parameters for $\mathcal{D}_{\text{f}}$ compared to $\mathcal{D}_{\text{tr}}$ in the model via the Fisher Information Matrix and dampens these parameters to induce unlearning. We replace SSD in our benchmarks with Potion (Schoepf et al., 2024b) which builds upon SSD and is the SOTA in poison unlearning as defined by Goel et al. (2024) (see Fig. 1). Potion iteratively increases the number of parameters in the model that it modifies analogously to SSD until unlearning of the poison trigger occurs. Potion stops unlearning as soon as it detects that the accuracy of the model on the forget set drops below a preset threshold, indicating that the poison has been unlearned. Fine-tuning continues to train the model on only $\mathcal{D}_{\text{r}}$, relying on catastrophic forgetting to remove $\mathcal{D}_{\text{f}}$. Retraining from scratch trains a new model using only $\mathcal{D}_{\text{tr}}\setminus\mathcal{D}_{\text{f}}$. Ascent performs gradient ascent on $\mathcal{D}_{\text{f}}$. NPO (Zhang et al., 2024a) is an alignment-inspired loss function for unlearning that overcomes the catastrophic collapses of Gradient Ascent. This is achieved via a reference-model based loss calculation.
@@
Investigation of unlearning limitations. Goel et al. (2024) studies the effect of partial discovery on unlearning method effectiveness as shown in Fig. 1. For unlearning poisoned data, they demonstrate that all methods apart from Foster et al. (2024) fail to remove the effect of the poison when only a subset of the poisoned examples is given (partial discovery). They also consider another setup, inter-class confusion, where samples from two classes are swapped with each other, where no method shows satisfactory performance in the partial discovery case. Zhao et al. (2024) identify interpretable characteristics of forget sets that substantially affect the difficulty of the problem, but they don’t consider the statistical regularity of the forget set in their investigation. Rezaei et al. (2025) show that standard unlearning methods designed for erasing knowledge may fail at restoring knowledge present prior to corruption. This relates to our metrics of healing, where we are interested in redirection to the correct label, rather than just avoiding predicting the incorrect (e.g. poisoned) label. However, their investigation is in the context of LLMs and they don’t establish a distinction between tasks of different discovery rates or regularities.
4 Taxonomy of Unlearning Tasks and Method Limitations
@@
Taxonomy. Our first contribution is to propose a new taxonomy of unlearning tasks according to two dimensions: the discovery rate and the statistical regularity of the corrupted data that we wish to unlearn. The discovery-rate dimension was previously discussed and empirically investigated in Goel et al. (2024) but is insufficient to explain model failures, as methods such as Potion (Schoepf et al., 2024b) excel at some low discovery tasks but then fail at other full-discovery tasks. We further identify regularity as a key dimension that affects the behaviour of unlearning algorithms both on its own and in conjunction with the discovery rate. Our key observation is that, as shown in Fig. 1, each prior method may excel on a slice of our 2D space of unlearning tasks but fails catastrophically everywhere else. We discuss the limitations of different algorithms below.
@@
Discovery rate. Unlearning methods that are designed for the traditional unlearning setting of full discovery fail catastrophically for lower discovery percentages as described by Goel et al. (2024). Retraining from scratch and unlearning methods that utilize the retain set, e.g. SCRUB, fail due to requiring a clean retain set for utility preservation, but in the partial discovery case, the retain set is contaminated and thus leads to reintroduction of corruptions (see Fig. 1 (b, e)). Methods such as Potion and SSD that only use $\mathcal{D}_{\text{f}}$ cause significant model utility damage due to the missing repair step on $\mathcal{D}_{\text{r}}$ (Goel et al., 2024; Foster et al., 2024).
@@
Regularity refers to the self-similarity of the corrupted data $\mathcal{D}_{\text{c}}$. Informally, we say that a corruption has low regularity if the corrupted data does not share any structure or common patterns. For example, a random mislabelling of a randomly-chosen subset of examples is a low-regularity corruption because the self-similarity between samples that end up with the same label change is low. On the other hand, a poisoning attack that introduces a trigger to each image whose label it wants to redirect results in a high-regularity corruption since all corrupted images share the visual pattern of the poison trigger and are all labeled in the same way. In the context of the commonly studied problem of full discovery unlearning, class unlearning (Golatkar et al., 2020) and random subset unlearning (Chundawat et al., 2023) are high and low regularity tasks, respectively. Several measures could be used to formalize regularity. We advocate for the consistency score (C-score) of Jiang et al. (2020), that measures the expected accuracy of a model on held-out samples from the training set. Their precomputed CIFAR10 scores align with the regularity rankings of our tasks (see Fig. 13 in Jiang et al. (2020)). Adapted to $\mathcal{D}_{\text{c}}$, C-scores can be computed with $n=0,1,...|\mathcal{D}_{\text{c}}|$ for instances $x$ and labels $y$, where $\mathbb{E}^{r}$ is empirical averaging over $r$ samples (Jiang et al., 2020):
[…]
@@
Experimental setup. We evaluate tasks of low, medium, and high regularity with ten discovery rates for each task (10%-100%). Low regularity is represented by random label swaps of $n$ samples to an incorrect label (Maini et al., 2023). In this setting, there is no regularity linking the samples or the new labels together. Medium regularity uses the inter-class confusion setup of Goel et al. (2022) where $n$ samples of two classes are swapped with each other. The highest regularity task uses a poison trigger to redirect to class 0 as defined by Goel et al. (2024). We measure the accuracy of $\theta_{u}$ on two types of data: (i) corrupted data (using the clean labels to compute accuracy), and (ii) non-corrupted data. The former is used to measure “healing”, i.e. whether the label prediction has been “redirected” to the correct label, while the latter is used to measure “utility”, i.e. ability to predict correctly on non-corrupted data. We perform experiments on an A100 (40GB) GPU with ResNet-9 (He et al., 2016) and Vision Transformer (ViT) (Dosovitskiy et al., 2021) models, stochastic gradient descent (SGD) and Adam (Kingma and Ba, 2014) optimizers, and CIFAR10 (Krizhevsky et al., 2009) and Street View House Numbers (SVHN) (Netzer et al., 2011) to show generalizability across models, optimizers and datasets. GitHub: https://github.com/google-deepmind/rem
5 Redirection for Erasing Memory (REM)
We now introduce our core contribution, Redirection for Erasing Memory (REM), the first unlearning method that is performant across our 2D space spanning different regularities and discovery rates, without requiring specialized hyperparameter tuning for each region. The design of REM is based on the lessons discussed in the previous section: (i) In the setting of partial discovery, finetuning on $\mathcal{D}_{\text{r}}$ causes reinforcing or reintroducing corruption (especially for high-regularity corruptions), but not using $\mathcal{D}_{\text{r}}$ at all would lead to low model utility; (ii) ETD can eliminate the effect of low-regularity corruptions by simply dropping out the memorization partition, but this operation will not remove the effect of high-regularity corruptions, as those are encoded in the generalization partition.
Let $\theta_{o_{1}}=\mathcal{A}(\mathcal{D}_{\text{tr}})$ denote the parameters of the original model, trained using $\mathcal{A}$ on $\mathcal{D}_{\text{tr}}$. Unlike ETD, REM’s choice of $\mathcal{A}$ can be any standard training procedure. REM then applies a post-processing $\mathcal{U}$ on top of $\theta_{o_{1}}$ to unlearn corrupted data by redirecting the effect of corruptions to newly-added parameters $\theta_{o_{2}}$ that are then dropped. We emphasize that, while this design draws inspiration from ETD, it is a fundamentally distinct methodology: while ETD attempts at training time to partition the network into generalization and memorization, REM attempts at post-processing time to partition the network into parameters that are unaffected by the corruptions and ones that capture all corruptions. Furthermore, REM adopts a unique masking strategy to capture high regularity corruptions that cannot be captured by ETD. We detail each REM step from Fig. 2 below and in Alg 1.